Solve each system of equations.

6x + 5y = 10
7x – 3y = 24

Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

6x + 5y = 10
7x – 3y = 24
-----
Multiply top equation by 3
Multiply bottom equation by 5
------
18x + 15y = 30
35x - 15y = 120
----
Add and solve for "x")
53x = 150
x = 150/53
-----------------
Solve for "y" using 6x + 5y = 10
-----
6(150/53) + 5y = 10
900/53 + 5y = 10
----
900/53 + 5y = 530/53
------
5y = -370/53
y = -370/265
-----------------
Cheers,
Stan H.
================

Answer by checkley79(3341)

6x + 5y = 10 multiply by 3.
7x – 3y = 24 multiply by 5 & add.
18x+15y=30
35x-15y=120 add.
-----------------
53x=150
x=150/53
x=2.83 ans.
6*2.83+5y=10
16.98+5y=10
5y=10-16.98
5y=-6.98
y=-6.98/5
y=-1.396 ans.
Proof:
7*2.83-3*-1.396=24
19.81+4.17=24
23.98=24